OpenStudy (anonymous):

What is the value of the function y = 0.5x – 10 when x = 9?

OpenStudy (anonymous):

plug in 9 and multiply it out?

OpenStudy (anonymous):

y=0.5*9-10=4.5-10=-5.5

OpenStudy (anonymous):

0.5*9 - 10

OpenStudy (anonymous):

-5.5 answer

OpenStudy (anonymous):

@dtuniyants not to be mean or anything, but please provide an explanation since "direct answers can't be given". trust me i have had my fair share of being told this. don't want to get you into trouble

OpenStudy (anonymous):

just plug in the value of the variable into its place and solve

OpenStudy (anonymous):

there we go.